# Simultaneous Polarization Conversion and Anomalous Reflection with Anisotropic Printed-Circuit-Board (PCB) Metagratings

### v1

@@ II.1 Formulation
$\begin{split}&k_{t_{m}}\!\overset{\triangle}{=}\!k\left(m\lambda/\Lambda+\sin\theta_{\mathrm{in}}\right)\!\overset{\triangle}{=}\!k_{i}\sin\theta_{m,i}\\ &\beta_{m,i}\!=\!\sqrt{k_{i}^{2}-k_{t_{m}}^{2}}\overset{\triangle}{=}k_{i}\cos\theta_{m,i},\ \mathfrak{Im}\left\{\beta_{m,i}\right\}\!<\!0,\end{split}$ (14)
@@ II.2 Simultaneous polarization conversion
To attain perfect anomalous reflection, i.e. redirect all the incident power from $\theta_{\mathrm{in}}$ towards the anomalous $\theta_{\mathrm{out}}$ angle, we first wish to prevent any power from being coupled to the fundamental ($0^{\mathrm{th}}$ order) TE or TM modes in Eq. (15). This is done by demanding destructive interference between the incident wave’s specular reflection off the PEC-backed dielectric and the $0^{\mathrm{th}}$ order FB-mode generated by the MG, which are associated, respectively, with the right-hand terms in Eqs. (2) and (9) for $m\!=\!0$. Altogether, formulating this destructive interference requirement for the two polarizations independently provides two equations requiring a specific relation between the induced dipole moments per unit length $\tilde{I}_{1}$ and $\tilde{I}_{2}$ to the MA locations and orientation DOFs ($h_{1},h_{2},\psi_{1},\psi_{2}$ and $d$). These relations are given by
@@ III.1 Perfect TE-to-TM polarization conversion
To verify the general theoretical scheme developed in Section II.2, we apply it to design a PCB-based MG, aimed at reflecting an illuminating TE-polarized plane wave impinging upon it at a $\theta_{\mathrm{in}}\!=\!10^{\circ}$ incidence angle, towards an anomalous $\theta_{\mathrm{out}}\!=\!-60^{\circ}$ trajectory, whilst simultaneously converting its polarization to TM ($E^{\mathrm{TE}}_{\mathrm{out}}\!=\!0$) at the output. Note, that due to the reciprocity theorem, this MG also operates as a TM-to-TE polarization converter when swapping $\theta_{\mathrm{in}}$ and $\theta_{\mathrm{out}}$, respectively.
@@
Figure 3: Symmetric private case of the (general) two-dipole scatterers per unit cell MG, featured in Fig. 1. Herein, the two identically scaled meta-atoms are equidistant from the PEC ($h_{1}\!=\!h_{2}$) and are both located on the dielectric-air interface; they are inversely rotated with respect to one another ($\psi_{1}=-\psi_{2}$), featuring mirror symmetry within each unit cell. This symmetric MG configuration is optimally compatible for TE-to-TM (or TM-to-TE) polarization conversion, simultaneously with anomalous reflection.
Next, we validate the performance of this solution using a full-wave simulation in CST Microwave Studio. The MAs are implemented as realistic passive and lossy copper dog-bone shaped polarizable particles, positioned at the prescribed coordinates $(y,z)=(0,-h)$ and $(y,z)=(d,-h)$, $h\!\overset{\triangle}{=}\!h_{1}\!=\!h_{2}$, and tilted by $\psi\!\overset{\triangle}{=}\!\psi_{1}\!=\!-\psi_{2}$, with the precise values featured above, found by theory. Their (conductor) width and thickness are set to $w\!=\!5\mathrm{mil}=127\mu\mathrm{m}$ and $t\!=\!18\mu\mathrm{m}$, respectively, compatible with standard PCB laminates and fabrication capabilities [Fig. 3]. By virtue of to the MAs position and orientation symmetry and the received requirement for identical current amplitudes, it follows that their required physical sizes should also be identical, as elaborated in Section III.2. A brief parametric sweep is performed to find these optimal MA lengths, and relating to the dimensions in Fig. 3, these read $(\ell_{1},\ell_{2},\ell_{3})=(0.206,0.061,0.056)\lambda$.
Figure 4: Field distributions $\Re\{E_{x}^{\mathrm{tot}}(y,z)\}$ (left, TE polarization) and $\Re\{E_{y}^{\mathrm{tot}}(y,z)\}$ (right, TM polarization) across a single $y$-period, corresponding to the MG from section III.1; designed to convert a TE polarized plane wave impinging upon the MG at $\theta_{\mathrm{in}}\!=\!10^{\circ}$, into a TM polarized plane wave anomalously reflected towards $\theta_{\mathrm{out}}\!=\!-\!60^{\circ}$. The black arrows indicate the direction of propagation, and the dashed white circles mark the two meta-atoms situated at $(y,z)\!=\!(\pm d/2,-h)$. Analytical predictions (a) following Eq. (III.1) are compared with full-wave simulation of passive copper dog-bone shaped polarizable meta-atoms (b).
@@ III.3 Experiment
After providing analytical insights regarding the MG design presented in Section III.1, we turn to verify experimentally the proposed concept. To this end, we strive to fabricate a prototype of this TE-to-TM $10^{\circ}$-to-$-60^{\circ}$ MG case study, building upon the corresponding analytical and full-wave simulation results [Fig. 4]. Considering now a more realistic version of the Rogers RO4350B laminate, featuring loss tangent of $\mathrm{tan}\delta\!=\!0.0037$ and an inherent anisotropic (uniaxial) permittivity of $\varepsilon_{zz}\!=\!3.66\varepsilon_{0}$ in the direction normal to the substrate and $\varepsilon_{xx}\!=\varepsilon_{yy}\!=\!3.75\varepsilon_{0}$ in the laminate plane [54], the optimal working point of the MG is slightly modified. While we retain the same MG thickness $h$ and MA tilt angle $\psi$ that were set by the semianlaytical design scheme for operation at $20$ GHz (Sections III.1 and III.2), the MA dimensions leading to the highest performance vary slightly (by ca. $2\%$) with respect to the ones found in Section III.1 due to this laminate anisotropy and loss, reading $(\ell_{1},\ell_{2},\ell_{3})=(0.201,0.06,0.055)\lambda$. Taking into account the additional dissipation in the dielectric somewhat reduces the overall cross-polar anomalous reflection efficiency, rating at $\eta_{c}\!=\!90.27\%$, with $0.98\%$ of the incident power being coupled to the residual modes and all other $8.75\%$ being absorbed in the copper traces and in the lossy substrate.
@@
This fabrication-ready MG layout was realized as a $9^{\prime\prime}\!\times\!12^{\prime\prime}$ board prototype using standard PCB manufacturing techniques, and is shown in Fig. 5(a). Due to limited availability of commercial laminate sizes and fabrication tolerances, the final thickness of the manufactured board was $77.5\mathrm{mil}$ instead of the required $80.5\mathrm{mil}$.
@@
With this regard, the experimentally measured and full-wave simulated reflected efficiencies, corresponding to the prototype MG at the actual experimental conditions, are presented in Fig. 7. Therein, the coupling efficiency is evaluated using
@@
comparing the peak measured (TM) gains of the MG $G_{\mathrm{MG}}(\theta_{\mathrm{out}})$ with the reference incident (TE) power in its absence $G_{\mathrm{ref}}(\theta_{\mathrm{in}})$, as in [33]. In reasonable agreement with the simulation predicted $\eta_{c}\!=\!90.23\%$, the experiment measured coupling efficiency peaking at the central frequency of $20.1\mathrm{GHz}$ rates at a high $\eta_{c}\!=\!85.9\%$, with a deviation of merely $4.33\%$. The received frequency response further validates the correspondence between the theoretical model and the recorded results.

### v2

@@ II.1 Formulation
$\begin{split}&k_{t_{m}}\!\overset{\triangle}{=}\!k_{1}\left(m\lambda/\Lambda+\sin\theta_{\mathrm{in}}\right)\!\overset{\triangle}{=}\!k_{i}\sin\theta_{m,i}\\ &\beta_{m,i}\!=\!\sqrt{k_{i}^{2}-k_{t_{m}}^{2}}\overset{\triangle}{=}k_{i}\cos\theta_{m,i},\ \mathfrak{Im}\left\{\beta_{m,i}\right\}\!<\!0,\end{split}$ (14)
@@ II.2 Simultaneous polarization conversion
To attain perfect anomalous reflection, i.e. redirect all the incident power from $\theta_{\mathrm{in}}$ towards the anomalous $\theta_{\mathrm{out}}$ angle, we first wish to prevent any power from being coupled to the fundamental ($0^{\mathrm{th}}$ order) TE or TM modes in Eq. (15)  [^2] Note that due to the anisotropic nature of the rotated MAs, the incident field generally induces dipoles with non-vanishing projections on both the $x$ and $y$ axes, namely, generating scattered FB modes of both TE and TM polarizations. This is true even if the incident wave is strictly single-polarized (i.e., even if $E_{\mathrm{in}}^{\mathrm{TE}}=0$ or $E_{\mathrm{in}}^{\mathrm{TM}}=0$), as shall be demonstrated in Section III.. This is done by demanding destructive interference between the incident wave’s specular reflection off the PEC-backed dielectric and the $0^{\mathrm{th}}$ order FB-mode generated by the MG, which are associated, respectively, with the right-hand terms in Eqs. (2) and (9) for $m\!=\!0$. Altogether, formulating this destructive interference requirement for the two polarizations independently provides two equations requiring a specific relation between the induced dipole moments per unit length $\tilde{I}_{1}$ and $\tilde{I}_{2}$ to the MA locations and orientation DOFs ($h_{1},h_{2},\psi_{1},\psi_{2}$ and $d$). These relations are given by
@@ III.1 Perfect TE-to-TM polarization conversion
To verify the general theoretical scheme developed in Section II.2, we apply it to design a PCB-based MG, aimed at reflecting an illuminating TE-polarized plane wave impinging upon it at a $\theta_{\mathrm{in}}\!=\!10^{\circ}$ incidence angle, towards an anomalous $\theta_{\mathrm{out}}\!=\!-60^{\circ}$ trajectory, whilst simultaneously converting its polarization to TM ($E^{\mathrm{TE}}_{\mathrm{out}}\!=\!0$) at the output. As laid out in Section II, the anisotropic rotated MAs incorporated to introduce coupling between the TE and TM field polarizations may give rise to both TE and TM propagating FB modes in all the relevant (allowed according to the FB theorem) set of angles – even if the incident field only features TE-polarized components. In particular, in the case considered herein, these propagating modes correspond to two (i.e., TE and TM) waves in the specular reflection direction ($m=0$ mode), and two (i.e., TE and TM) waves in the anomalous reflection trajectory ($m=-1$). Correspondingly, the synthesis goal would be to suppress the three spurious modes and direct all of the incident TE power towards the anomalous TM mode. Specifically, this implies utilization of the devised model with $\eta_{-1}^{\mathrm{TE}}=0$ in Eq. (18).
@@
Figure 3: Symmetric private case of the (general) two-dipole scatterers per unit cell MG featured in Fig. 1, corresponding to the configuration synthesized and analyzed in Section III. Herein, as opposed to Fig. 1(b), the two identically scaled meta-atoms are equidistant from the PEC ($h_{1}\!=\!h_{2}$) and are both located on the dielectric-air interface; they are inversely rotated with respect to one another ($\psi_{1}=-\psi_{2}$), featuring mirror symmetry within each unit cell.
Next, we validate the performance of this solution using a full-wave simulation in CST Microwave Studio. The MAs are implemented as realistic passive and lossy copper dog-bone shaped polarizable particles, positioned at the prescribed coordinates $(y,z)=(0,-h)$ and $(y,z)=(d,-h)$, $h\!\overset{\triangle}{=}\!h_{1}\!=\!h_{2}$, and tilted by $\psi\!\overset{\triangle}{=}\!\psi_{1}\!=\!-\psi_{2}$, with the precise values featured above, found by theory. Their (conductor) width and thickness are set to $w\!=\!5\mathrm{mil}=127\mu\mathrm{m}$ and $t\!=\!18\mu\mathrm{m}$ respectively [Fig. 3], chosen as to comply well with conventional PCB manufacturing resolution tolerances and standard (commercially available) electro-deposited copper foils. By virtue of the MAs position and orientation symmetry and the received requirement for identical current amplitudes, it follows that their required physical sizes should also be identical, as elaborated in Section III.2.
To obtain optimal MA lengths $(l_{a},l_{b},l_{c})$ relating to the dimensions in Fig. 3 – namely, the geometry that would yield an effective polarizability which would support the required induced currents once excited by the incident fields – an initial heuristic setting for their proportions was adopted, $(l_{a,i},l_{b,i},l_{c,i})=(0.245,0.073,0.067)\lambda$. Subsequently, these were multiplied by a common scaling factor $\xi$ over which we performed a brief parametric sweep in CST under periodic boundary conditions to find the (optimal) lengths, seeking the highest coupling efficiency of the TE-polarized incident wave to the anomalous $m\!=\!-1$ TM-polarized FB mode. Executing this procedure for the MG parameters and functionality specified in this section resulted in such a valid solution ($\xi\!=\!0.84$), setting the MA dimensions to $(l_{a},l_{b},l_{c})=(0.206,0.061,0.056)\lambda$.
Figure 4: Field distributions $\Re\{E_{x}^{\mathrm{tot}}(y,z)\}$ (left, TE polarization) and $\Re\{E_{y}^{\mathrm{tot}}(y,z)\}$ (right, TM polarization) across a single $y$-period, corresponding to the MG from Section III.1; designed to convert a TE polarized plane wave impinging upon the MG at $\theta_{\mathrm{in}}\!=\!10^{\circ}$, into a TM polarized plane wave anomalously reflected towards $\theta_{\mathrm{out}}\!=\!-\!60^{\circ}$. The white arrows indicate the direction of propagation, and the dashed white circles mark the two meta-atoms situated at $(y,z)\!=\!(\pm d/2,-h)$. Analytical predictions (a) following Eq. (III.1) are compared with full-wave simulation of passive copper dog-bone shaped polarizable meta-atoms (b).
@@
We should emphasize that while in this section we demonstrated and verified our methodology for a single case study, with TE-to-TM polarization conversion alongside anomalous reflection from $\theta_{\mathrm{in}}\!=\!10^{\circ}$ to $\theta_{\mathrm{out}}\!=\!-60^{\circ}$, the scheme developed in Section II is entirely universal, and may effectively be employed for a broad range of input-output angle pairs, including shallow ones, as well as different incident wave polarization states. To further illustrate this universality, another case study is included in Appendix D, wherein a TE-polarized field from $\theta_{\mathrm{in}}\!=\!70^{\circ}$ is converted into a TM-polarized wave reflected towards $\theta_{\mathrm{out}}\!=\!-5^{\circ}$ with high efficiency, demonstrating an even more challenging (in terms of deflection angles) functionality. As also denoted in the beginning of this subsection, by virtue of the reciprocity theorem [4], these results imply the effectiveness of the devised technique also for TM-polarized incident fields, impinging upon the MG from either near-broadside or near-grazing angles.
Lastly, it is worthy of note that while the theoretical derivation guarantees that there should exist a passive and lossless scatterer which could support the prescribed functionality, it does not directly specify the geometry of such a MA. Consequently, although the representative cases considered in Section III and Appendix D (as well as others not shown) do demonstrate the substantial versatility of the proposed concept and synthesis methodology, it is difficult to assess accurately the limitations of the devised scheme in the context of practical realization in the microwave regime. A potential constraint is related to the lateral dimensions of the MA, which must remain deeply subwavelength to allow theoretical consideration as a rotated dipole line (i.e., having a dominant dipole moment), necessary to facilitate our analytical model and corresponding synthesis method (Section II.2). When large induced currents are required (e.g., when considering near-grazing angle of incidence/reflection, $|\theta_{\mathrm{in/out}}|\!\gtrsim\!75^{\circ}$), the MA polarizability should be significant as well, which generally requires larger dipolar scatterers (larger $l_{a}$). This may become a limiting factor in case one fails to sufficiently tune the resonant conditions via the loading arms ($l_{b},l_{c}$). While a more rigorous treatment in the MA physical design is beyond the scope of this paper, such an extension, e.g., relying on [4], [36], will be incorporated into the methodology in future work.
@@ III.3 Experiment
After providing analytical insights regarding the MG design presented in Section III.1, we turn to verify experimentally the proposed concept. To this end, we strive to fabricate a prototype of this TE-to-TM $10^{\circ}$-to-$-60^{\circ}$ MG case study, building upon the corresponding analytical and full-wave simulation results [Fig. 4]. Considering now a more realistic version of the Rogers RO4350B laminate, featuring loss tangent of $\mathrm{tan}\delta\!=\!0.0037$ and an inherent anisotropic (uniaxial) permittivity of $\varepsilon_{zz}\!=\!3.66\varepsilon_{0}$ in the direction normal to the substrate and $\varepsilon_{xx}\!=\varepsilon_{yy}\!=\!3.75\varepsilon_{0}$ in the laminate plane [55], the optimal working point of the MG is slightly modified. While we retain the same MG thickness $h$ and MA tilt angle $\psi$ that were set by the semianlaytical design scheme for operation at $20$ GHz (Sections III.1 and III.2), the MA dimensions leading to the highest performance vary slightly (by ca. $2\%$) with respect to the ones found in Section III.1 due to this laminate anisotropy and loss, reading $(l_{a},l_{b},l_{c})=(0.201,0.06,0.055)\lambda$. Taking into account the additional dissipation in the dielectric somewhat reduces the overall cross-polar anomalous reflection efficiency, rating at $\eta_{c}\!=\!90.27\%$, with $0.98\%$ of the incident power being coupled to the residual modes and all other $8.75\%$ being absorbed in the copper traces and in the lossy substrate  [^5] Note, that the simulated $\eta_{c}$ is the value of the relevant entry in the scattering matrix of the devised configuration, defined and simulated in CST. This scattering parameter, quantifying the scattered TM power coupled to the desired TM $m=-1$ FB mode relative to the overall incident TE power, is a part of the standard output by CST within such a Floquet port simulation..
@@
This fabrication-ready MG layout was realized as a $9^{\prime\prime}\!\times\!12^{\prime\prime}$ ($22.86\mathrm{cm}\!\times\!30.48\mathrm{cm}$) board prototype using standard PCB manufacturing techniques, and is shown in Fig. 5(a). Due to limited availability of commercial laminate sizes and fabrication tolerances, the final thickness of the manufactured board was $77.5\mathrm{mil}$ ($1.969\mathrm{mm}$) instead of the required $80.5\mathrm{mil}$ ($2.0446\mathrm{mm}$).
@@
With this regard, the experimentally measured and full-wave simulated reflected efficiencies, corresponding to the prototype MG at the actual experimental conditions, are presented in Fig. 7. Therein, the coupling efficiency is evaluated using [33]
@@
comparing the peak measured (TM) gain of the MG, $G_{\mathrm{MG}}(\theta_{\mathrm{out}})$, with the reference incident (TE) power in its absence, $G_{\mathrm{ref}}(\theta_{\mathrm{in}})$, evaluated directly from the measured (gain) radiation pattern [54]. As discussed in [33], the form of Eq. 34 allows quantification of the power carried by the various scattered beams (corresponding to the various FB modes) for a given finite aperture, properly calibrating for the different effective aperture lengths for different observation angles via the factor ${\cos\theta_{\mathrm{in}}}/{\cos\theta_{\mathrm{out}}}$. In reasonable agreement with the simulation predicted $\eta_{c}\!=\!90.23\%$, the experiment measured coupling efficiency peaking at the central frequency of $20.1\mathrm{GHz}$ rates at a high $\eta_{c}\!=\!85.9\%$, with a deviation of merely $4.33\%$. The received frequency response further validates the correspondence between the theoretical model and the recorded results.
@@ IV Conclusion
Overall, the work fills critical gaps which currently exist in MS and MG analysis and synthesis. Regarding the latter, it proposes an architecture and lays out a systematic framework for extending the wavefront and polarization manipulation capabilities beyond single-polarized TE [8, 9, 15] or TM [10, 35, 36] MGs, allowing simultaneous polarization conversion and beam deflection with high efficiency in an appealing PCB platform  [^6] Indeed, the experimentally evaluated efficiency of $85.9\%$ for the anomalous reflection and polarization conversion MG devised herein is comparable with other measured beam manipulating MG efficiencies of $\sim 90\%$ reported in the past [18].. In the context of the former, it avoids the homogenization approximation hurdles via the formulated semianalytical model allowing swift effective practical designs for realizing this challenging functionality. Correspondingly, while some MSs implementing combined wave trajectory and polarization conversion have been theoretically envisioned before  [^7] Being only theoretical and neglecting loss, it is somewhat difficult to compare the solutions in [44, 45] to the designs presented herein. If losses are neglected for the design in Section III.1, the efficiency reaches $99.4\%$, almost identical to the optimal unitary MS solutions. For the simulated designs herein and in [45], one may also comment on the $90\%$ fractional bandwidth (with respect to the peak efficiency), finding a bandwidth of about $5\%$ for the results presented in Fig. 7 of Section III.3 compared to about $2\%$ in [45]., the non-trivial need to translate the homogenized abstract GSTC-based highly anisotropic solutions into actual manufacturable layouts of numerous dense meta-atoms may require extensive full-wave optimization [45] or end up with only a conceptual solution at hand [44]. In contrast, the MG approach devised herein directly yields fabrication-ready devices, with sparse element distributions that may also reduce manufacturing complexity. Importantly, utilizing the minimally required number of degrees of freedom in conjunction with the analytical model facilitates valuable analytical insights (Section III.2), which shed light also on the symmetries apparent in related MS designs [45], in which such observations are sometimes more challenging to achieve.
@@ Appendix C Perfect TM-to-TE polarization conversion and anomalous reflection
Appendix D Wide-angle simultaneous polarization conversion and anomalous reflection with shallow-angle incidence
To further exemplify the versatility of our synthesis framework, we showcase an additional second case study to the one presented in Section III.1, considering both a relatively shallower incidence angle and a more acute output angle. Herein, a TE-polarized wave impinging at a $\theta_{\mathrm{in}}\!=\!70^{\circ}$ incidence angle, is reflected as a TM-polarized wave, exiting at an anomalous $\theta_{\mathrm{out}}\!=\!-5^{\circ}$ output angle. The analytically-derived DOFs values for this specific case read: $\tilde{I}_{1}\!=\!(0.0487\!+\!j0.0378)E^{\mathrm{TE}}_{\mathrm{in}}\Lambda/\eta_{1}$, $\tilde{I}_{2}\!=\!(\!-\!0.037\!-\!j0.0493)E^{\mathrm{TE}}_{\mathrm{in}}\Lambda/\eta_{1}$, $d\!=\!\Lambda/2\!=\!7.2088\mathrm{mm}$, $\psi_{1}\!=\!-\!\psi_{2}\!=\!1.0305\ \mathrm{rad}$ and $h_{1}\!=\!h_{2}\!=\!0.1428\lambda\!=\!2.1403\mathrm{mm}$.
The theoretically predicted $100\%$ perfect coupling efficiency, is met with a high $\eta_{c}\!=\!93.29\%$ received from a full-wave simulation including realistic copper MAs and a lossy and anisotropic RO4350B substrate, as used in Section III.1. Moreover, merely $0.73\%$ of the incident power is coupled to other spurious modes, implying that the residual $5.98\%$ is absorbed as loss by the device. Note, that optimal MA dimensions were again found through a similar parametric sweep procedure as described in Section III.1, resulting in $(l_{a},l_{b},l_{c})=(0.164,0.049,0.048)\lambda$; all other parameters remained unchanged, as listed for the simulation in Section III.1. In analogy to Fig. 4, Fig. 8 showcases the fields distributions visualizing the good correspondence between theory and simulation, once more verifying our semianalytical synthesis scheme.
Figure 8: Field distributions $\Re\{E_{x}^{\mathrm{tot}}(y,z)\}$ (left, TE polarization) and $\Re\{E_{y}^{\mathrm{tot}}(y,z)\}$ (right, TM polarization) across a single $y$-period, corresponding to the MG from Appendix D; designed to convert a TE polarized plane wave impinging upon the MG at $\theta_{\mathrm{in}}\!=\!70^{\circ}$, into a TM polarized plane wave anomalously reflected towards $\theta_{\mathrm{out}}\!=\!-\!5^{\circ}$. All other parameters are identical to the ones listed in the caption of Fig. 4.
We should note that due to the reciprocity theorem, the case studies presented in Section III.1 and Appendix D demonstrate, in fact, also the capability of the proposed scheme to devise TM-to-TE polarization converting anomalous reflectors, for both shallow and near-normal angles of incidence, respectively. In particular, reciprocity dictates that $\eta_{c}$ should remain identical when replacing the excitation and observation probes, i.e., when illuminating the MG by a TM-polarized field incident from what was defined in Section III.1 and Appendix D as $\theta_{\mathrm{out}}$ and considering the coupling towards $\theta_{\mathrm{in}}$. Indeed, when simulating the MG designed in Section III.1 (unchanged) with an illumination of a TM-polarized wave from $\theta_{\mathrm{in}}=60^{\circ}$, the coupling to $\theta_{\mathrm{out}}=-10^{\circ}$ as retrieved from the simulation is $96.8\%$, with $0.64\%$ coupled to spurious modes and $2.56\%$ dissipated as loss (vs. the original efficiency rates $96.5\%$, $0.68\%$, and $2.82\%$, respectively). Similarly, when simulating the MG designed in Appendix D (unchanged) with an illumination of a TM-polarized wave from $\theta_{\mathrm{in}}=5^{\circ}$, the coupling to $\theta_{\mathrm{out}}=-70^{\circ}$ as retrieved from the simulation is $93.53\%$, with $0.62\%$ coupled to spurious modes and $5.68\%$ dissipated as loss (vs. the original efficiency rates $93.29\%$, $0.72\%$, and $5.75\%$, respectively). The minor deviations of $\sim 0.05\%$ from the reciprocity theorem predictions are attributed to numerical errors within the full-wave solver.